OpenStudy (anonymous):

If (x+7)(x-2) = ax^2 + bx + c for all values of x, where a, b, and c are constants, what is the value of a + b - c

OpenStudy (amoodarya):

expand (x+7)(x-2) then compare rhs ,lhs

OpenStudy (usukidoll):

that's what I was thinking too... if we use the FOIL method to expand (x+7)(x-2) we can figure out what a, b, and c is

OpenStudy (amoodarya):

\[(x+7)(x-2)=x^2+5x-14=ax^2+bx+c\\ax^2=x^2\\bx=5x\\c=-14\]

OpenStudy (amoodarya):

another method is : put for example x=0,1,2 both side then you have a system of equation to find a,b,c

OpenStudy (anonymous):

ohhh okay! It makes much more sense now! Thanks guys!! :)

OpenStudy (usukidoll):

it looked like a simple proof problem where we start from the left side and have it equal to the right side, so by expanding (x+7)(x-2) using FOIL we will have \[x^2+5x-14 \]. Then we compare that equation to \[ax^2+bx+c \] which is the standard quadratic equation.

OpenStudy (anonymous):

Got it! Thank you!